-- HarryRenshall - 16 Jan 2008

Week of 080317

Open Actions from last week: Respond to RAL what are the LHC experiment requirements for local RB and WMS at RAL.

Monday:

See the weekly joint operations meeting minutes

Additional Material:

ATLAS (AdG): are shifting their announced schedule to continue T1-T1 debugging this week and perform throughput tests next week (after Easter). They would like to see a simple interface to the T1 Gocdb downtimes where all T1 are summarised on one page.

CMS (DB): have released prodagent 7.1 with improvements on cleanup at sites. Still having delays getting global run data into the CAF. In T1-T1 commissioning only RAL-ASGC is missing.

LHCb (RS): An LHCb software week this week so not much activity on the grid. Testing migration of T1D1 data from STORM to tape under the IBM Tivoli Storage Manager.

Tuesday:

elog review: No new reports.

Experiment report(s):

CMS (DB): This morning patches were applied to the CERN CMS Castor instance to improve the performance of their CAF data access.

ATLAS (SC): Two activities - setting up the SRM2 CERN Castor environment and functional tests T1-T1 with NL-T1 as target. Tests to IN2P3 were mildly successful (2 ddm problems found) while those to PIC will have to be repeated. Some channels are slow so they are trying to tune FTS configurations. The intention is to build a Wiki page of suitable configurations for the T1-T1 combinations.

Core services (CERN) report: FTS will move to the so-called RAL channel setup, which separates out experiments more, next Tuesday.

DB services (CERN) report:

Monitoring / dashboard report:

Release update:

Questions/comments from sites/experiments:

AOB: It was agreed to move the RAL query on what WMS resources are needed at a T1 to the next F2F meeting.

Wednesday

elog review:

Experiment report(s):

Core services (CERN) report:

DB services (CERN) report:

Monitoring / dashboard report:

Release update:

Questions/comments from sites/experiments:

AOB:

Thursday

elog review:

Experiment report(s):

Core services (CERN) report:

DB services (CERN) report:

Monitoring / dashboard report:

Release update:

Questions/comments from sites/experiments:

AOB:

Friday

elog review:

Experiment report(s):

Core services (CERN) report:

DB services (CERN) report:

Monitoring / dashboard report:

Release update:

Questions/comments from sites/experiments:

AOB:

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2008-03-18 - HarryRenshall
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback